Each croquet ball in a set has a mass of 0.50

kg. The green ball, traveling at 14.6 m/s,

strikes the blue ball, which is at rest.

Assuming that the balls slide on a frictionless

surface and all collisions are head-on, find

the final speed of the blue ball in each of the

following situations:

a) The green ball stops moving after it

strikes the blue ball.

Answer in units of m/s.

b) The green ball continues moving after the

collision at 2.2 m/s in the same direction.

Answer in units of m/s.

c) The green ball continues moving after the

collision at 0.4 m/s in the same direction.

Answer in units of m/s.
Answer #1

let the speed of blue ball in each case is vB

initial speed of green ball , uG = 14.6 m/s

A)

final speed of green ball , vG = 0 m/s

Using conservation of momentum ,

m * uG = m * vG + m* vB

14.6 = 0 + vB

final speed of blue ball is 14.6 m/s

B)

final speed of green ball , vG = 2.2 m/s

Using conservation of momentum ,

m * uG = m * vG + m* vB

14.6 = 2.2 + vB

vB = 12.4 m/s

final speed of blue ball is 12.4 m/s

C)

final speed of green ball , vG = 0.4 m/s

Using conservation of momentum ,

m * uG = m * vG + m* vB

14.6 = 0.4 + vB

vB = 14.2 m/s

the final speed of blue ball is 14.2 m/s